OpenStudy (anonymous):

pls help will give medal stuck on math exam Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 The solution is where the points meet and I got .5,-4.5 but my answer choices are: A (1, 10) B (-1, 10) C (10, 1) D (1, -10)

OpenStudy (the_fizicx99):

Wait, I was also stuck on this. What does " less than or greater to " I think you mean less than or equal to.

OpenStudy (the_fizicx99):

Either this is horribly written or I also don't understand it. ._. @TuringTest

OpenStudy (turingtest):

draw the lines on a graph first....

OpenStudy (the_fizicx99):

Yea i did that, but what does less than or greater to mean?

OpenStudy (turingtest):

this is an exam for which class?

OpenStudy (the_fizicx99):

Math? >.>

OpenStudy (anonymous):

pls help i did mean greater than or equal to

OpenStudy (anonymous):

thats exactly how the question is written so idk man

OpenStudy (anonymous):

can you take a pic of the question? or are you in class? send it through here

OpenStudy (mathmale):

The way in which you've typed in this problem, DB123, is highly unusual. Are you able to use the equation editor (see the button below)? That's be so much clearer. I ask y ou to go back to the original question and make absolutely sure that you've copied the question exactly as it was presented by your teacher or whomever wrote the exam.

OpenStudy (mathmale):

Note that three people so far, including myself, question the way in which this problem is worded. That should tell you something. Personally, I'd ASSUME that the two inequalities are \[y \le x-5, y \le -x -4\] and graph y = x-5 and y=-x-4.

OpenStudy (the_fizicx99):

You got mathmale helping >.> my help won't be much of use here lol :3

OpenStudy (mathmale):

Personally, I'd graph y=x-5, then y=-x-4, on the same set of axes. Then I'd shade the area below each of these 2 lines (because of the " < " symbol). Then, one by one, I'd check each of the four given points. If any of them is within the area that you have shaded twice, then that one is a solution.

OpenStudy (mathmale):

Thank you very much, Luis!

OpenStudy (the_fizicx99):

xD

OpenStudy (mathmale):

:)

OpenStudy (mathmale):

@dumbo123: since you hadn't responded in quite some time, I checked and found that you were offline. I understand that emergencies do happen, and computers and Internet connections do fail now and then, but if you intentionally absented yourself, please inform your helpers in advance next time. I'd have enjoyed helping you arrive at a satisfactory answer to the question we've been discussing here.